OpenStudy (anonymous):

Debbie is borrowing 3 books from her friend who has a collection of 45 books. Which of the following expressions represents the number of ways Debbie can choose a group of 3 books?

OpenStudy (anonymous):

you are being asked how many ways to choose 3 out of a set of 40 this is called "forty choose 3" written as \[\dbinom{40}{3}\] and computed by \[\dbinom{40}{3}=\frac{40\times 39\times 39}{3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

ok replace 40 by 45 my mistake

OpenStudy (anonymous):

\[\dbinom{45}{3}=\frac{45\times 44\times 43}{3\times 2}\]